### v1

@@ Abstract
Hope is a complex and underexplored emotional state that plays a significant role in education, mental health, and social interaction. Unlike basic emotions, hope manifests in nuanced forms ranging from grounded optimism to exaggerated wishfulness or sarcasm, making it difficult for Natural Language Processing systems to detect accurately. This study introduces PolyHope V2, a multilingual, fine-grained hope-speech dataset comprising over 30,000 annotated tweets in English and Spanish. This resource distinguishes between four hope subtypes—Generalized, Realistic, Unrealistic, and Sarcastic—and enhances existing datasets by explicitly labeling sarcastic instances. We benchmark multiple pre-trained transformer models and compare them with large language models (LLMs) such as GPT-4 and Llama 3 under zero-shot and few-shot regimes. Our findings show that fine-tuned transformers outperform prompt-based LLMs, especially in distinguishing nuanced hope categories and sarcasm. Through qualitative analysis and confusion matrices, we highlight systematic challenges in separating closely related hope subtypes. The dataset and results provide a robust foundation for future emotion recognition tasks that demand greater semantic and contextual sensitivity across languages.
keywords:
Hope Speech Detection, Sarcasm Detection, Multilingual NLP, Emotion Recognition, Fine-grained Sentiment Analysis
@@ 1 Introduction
Recent improvements in Natural Language Processing (NLP) have enhanced applications in sentiment analysis, mental health assessments, social media monitoring, and educational platforms [1, 2, 3, 4, 5]. Despite recent progress, a persistent challenge in emotion recognition lies in identifying subtle and complex emotions, particularly hope, which is often overlooked in standard emotion taxonomies [6]. Unlike overt emotions such as anger, fear, sadness, or joy, hope is an inherently ambiguous emotion that exhibits differently across various contexts, and carries multiple emotional dimensions [7, 8]. Therefore, traditional NLP techniques struggle to capture subtle emotions and highlight the need for more refined, context-aware approaches to accurately detect nuanced emotional expressions, such as hope [9].
@@
Accurate detection of hope speech unlocks substantial benefits across real-world NLP applications, such as enhancing student engagement and success [10], understanding patients’ mental states [11, 12], and Non-medical prescription drug use (NPMDU). For example, within online education, instructors can leverage these systems as personalized support tools to accurately identify expressions of hope to better understand students’ motivations, interests, and emotional barriers—ultimately enabling targeted interventions [10]. Similarly, mental health professionals can leverage the natural progression of hope speech to better understand patients’ mental states, therapeutic outcomes, and concealed psychological factors over time. Research emphasizes that accurately interpreting psychological states requires identifying positive emotional expressions, particularly hope and optimism [11]. As a result, NLP systems significantly benefit from incorporating hope detection technologies, since integrating positive emotional indicators alongside negative sentiments enables more comprehensive, nuanced, and effective analytic models.
@@
Traditional NLP techniques struggle to detect hope due to its complex, dual nature as defined by Lazarus [13]. According to Lazarus, hope consists of two essential components: (i) a specific desire for positive outcomes, and (ii) an overall optimistic view of future events. The inherent contradictions within hope result in complicated textual interpretations, causing individual textual markers to be variably perceived as genuine goals, imaginary objectives, or sardonic remarks. Consequently, expressions intended to reflect hopeful thinking may inadvertently suggest agitation, depending on their textual context. For example, the sentence “I hope things finally change around here” reveals optimism but also hints at frustration, depending on how it is used in the text. Most emotion recognition systems trained on simplified affective categories (e.g., Ekman’s six basic emotions) tend to fail when processing hope, as a generic positive sentiment or as non-committal discourse, failing to grasp its subtleties [14, 7]. Thus, emotion recognition systems trained on basic emotion sets fail to correctly identify hope and often misclassify it either as a simple positive emotion or as a neutral, non-committal state.
@@
Additionally, recognizing hope is further complicated by sarcasm, as sarcastic expressions frequently employ positive language to communicate negative sentiments. The mismatch between literal meanings and intended emotions makes it challenging for machine learning and deep learning classifiers to distinguish genuinely hopeful statements from their sarcastic counterparts. For instance, sarcasm occurs when someone says, “Great, I really hope this mess gets worse,” even though the words themselves convey hope, but the semantic context reveals sarcastic negativity [15].
@@
The semantic incongruity between lexical content and emotional meaning makes emotion detection more challenging. The combination of figurative language and sarcasm complicates consistent emotion classification for computers since they primarily work with superficial word analysis instead of contextual interpretation [1]. Accurate context-aware emotion detection primarily depends on extensive, diverse datasets and sophisticated models [16]. However, there is a notable lack of annotated datasets explicitly related to hope, and existing datasets often either entirely omit hope or classify it merely as a subset of general positive emotions that cause incorrect predictions. Therefore, more comprehensive annotated datasets are necessary that explicitly capture hope through genuinely positive sentiments, unfounded expectations, and sarcastic expressions.
@@
In this study, we bridge the gap between emotion recognition and pragmatic language understanding and present a bilingual annotated hope speech dataset designed for hope detection in both English and Spanish. We explicitly annotate sarcasm within hope expressions and create rich annotation guidelines for different hope categories. We extend [9] hope categories (i) generalized hope, (ii) realistic hope, (iii) unrealistic hope or wishful thinking, and (iv) sarcasm. We develop sarcasm-resilient emotion classification models as a baseline to detect and evaluate the performance of the hope detection system in English and Spanish. This dataset can be used for advancing cross-linguistic emotion recognition research.
@@ 2 Literature Review
NLP has advanced beyond basic sentiment classification (positive, negative, or neutral) toward identifying more nuanced emotional states, including the subtle and underexplored emotion of hope [16, 6]. While emotions such as joy, fear, and regret have received substantial attention, hope remains comparatively underresearched, despite its significant role in mental health support [11], education [10], online communication, and social programs [17]. Hope consists of two distinct components that contribute to its complexity: (i) a general optimism inherent in individuals, and (ii) a specific expectation for positive outcomes [13]. Given its unique semantic, emotional, and contextual characteristics, recent research in hope speech detection has increasingly treated hope as an independent emotional category rather than merely grouping it under general positive emotions.
@@ 2.1 Methodologies for Hope Speech Detection
Different methodologies have been proposed for detecting hope speech and addressing challenges related to context, semantics, and multilingual content. Various studies have explored techniques to identify hope expressions within social media and textual documents [18, 19]. For instance, Balouchzahi et al. [9] introduced “PolyHope,” an NLP system that operates at semantic and contextual levels to detect nuanced expressions of hope. Similarly, another study [20] leveraged transformer-based models to differentiate between expressions of hope and regret, and highlighted complexities in jointly modeling these emotionally similar pairs. Despite these advancements, current methodologies lack in distinguishing hope expressions from sarcastic or superficially hopeful messages, particularly on social media platforms (e.g., X (previously Twitter), and Reddit). Deep learning models provide poor results because they often misinterpret pragmatic signals and contextual discourse markers [21]. These limitations underline the need to incorporate pragmatic and contextual awareness to improve model accuracy. Moreover, addressing these contextual challenges becomes even more critical when extending hope detection methodologies across multilingual and cross-cultural datasets.
@@
A growing body of work has explored multilingual and cross-lingual hope speech detection, recognizing its importance in linguistically diverse societies. Chakravarthi et al. [22, 23] introduced multilingual approaches to detect hope speech in English, Tamil, Malayalam, and Kannada. The authors explicitly emphasized equality and diversity and inclusion (EDI) to detect hope speech. Nevertheless, their models did not sufficiently account for distortions caused by sarcasm or unrealistic expectations embedded within hopeful expressions. Similarly, researchers such as Arunima et al. [24] have contributed to shared tasks like Hope Speech Detection for Equality, Diversity, and Inclusion (HopeEDI), which expanded datasets and evaluation benchmarks in multiple Indian languages. Additional efforts have targeted Urdu datasets, with studies exploring the transferability of hope speech models across languages and the need for language-specific semantic resources [25, 26]. Chakravarti [27] addressed hope speech detection in Tamil, English, and Malayalam, highlighting classification challenges due to code-mixed data. Moreover, Malik et al. [28] explored a joint multilingual and translation-based approach for hope speech detection in English and Russian using a fine-tuned Russian-RoBERTa model, achieving 94% accuracy and 80.24However, most of these models do not explicitly address the confounding effects of sarcasm, exaggeration, or unrealistic optimism expressed through hopeful language—factors that often distort detection accuracy in real-world applications. Moreover, research on low-resource and code-mixed language scenarios remains limited, although recent work has begun to tackle these challenges through cross-lingual transfer learning and zero-shot classification approaches [29].
@@ 2.2 Existing Datasets for Hope Speech Detection
Several datasets have been introduced to facilitate advancements in hope speech detection in various languages, such as Spanish [30, 31], Bengali [32]. For example, the “HopeEDI” dataset [22] introduced a multilingual corpus specifically targeting hopeful expressions related to equality, diversity, and inclusion (EDI). Additionally, the “IberLEF” Hope datasets [30, 31] extended hope speech detection into Spanish, and emphasized annotations around social causes and community identities. Another dataset, “BongHope” [32], presented a Bengali corpus for hope speech detection. Furthermore, the HOPE shared tasks at IberLEF provided comprehensive benchmarks and detailed insights into effective methodologies for annotating hope datasets [31, 33]. In parallel, the HASOC-Hope subtrack, part of the FIRE shared tasks, introduced additional multilingual datasets covering Hindi, Malayalam, and Tamil, further enriching the research landscape in this area [34]). Finally, the LREC Hope Speech dataset offers a multilingual resource tailored for cross-lingual and zero-shot hope speech detection, with a special focus on low-resource language scenarios [35].
@@ 2.3 Sarcasm in Hope Speech: A Complex Challenge
The detection of hope speech is highly impeded by the presence of sarcasm. Positive sentimental language through rhetorical expression functions as sarcasm to convey negative meanings while possessing a humorous or mocking quality. For example, “This one needs to go well as it happened previously”. This hopeful-sounding statement contains semantic skepticism through ironic speech. The model detection of hopeful speech can fail due to sarcastic statements that resemble hopeful statements. The improper evaluation of sentiment features occurs when basic analysis methods, such as positive polarity measurements, operate independently from detailed contextual analysis.
@@
The task of sarcasm detection is challenging because it requires contextual information, tonal cues, and background understanding [16]. Research on sarcasm detection within NLP has approached this problem in four ways: (i) Contextual embeddings (e.g., BERT, RoBERTa) [36], (ii) Multitask learning frameworks (to jointly learn emotion and sarcasm) [37], (iii) Contrastive learning approaches [38], and (iv) Discourse-aware models [39] that incorporate surrounding sentences or conversational history. Yet sarcasm-aware hope detection remains an unexplored intersection, and no existing hope detection dataset provides labeled sarcastic examples, as emphasized by Balouchzahi et al. [9] in their analysis of NPMDU-related textual data. To address this challenge, we annotate hope expressions along with sarcasm indicators and provide a richer, more realistic representation of hope in social discourse. Such annotation is critical for building models that can distinguish between sincere emotional optimism and sarcastic cynicism, as such a distinction is particularly important in domains like education, counseling, and mental health.
3 Dataset
In this work, we present PolyHope V2, a multilingual and fine-grained hope speech dataset consisting of around 30,000 annotated tweets in both English and Spanish. PolyHope V2 brings together the original English‐language PolyHope tweets [9], a newly harvested Spanish counterpart, and an additional sarcasm layer to create a multilingual, multi‑label resource for modelling hope on social media. Altogether the corpus contains 9,515 English and 20,442 Spanish tweets, each message annotated at the tweet level for hope, sarcasm, and the related fine‑grained categories described below. The next subsections outline how the English, Spanish, and sarcasm components were collected, cleaned, and aligned. The complete dataset is publicly available through the webpage [^1] https://www.codabench.org/competitions/5509/.
3.1 English
We relied on the PolyHope corpus [9] for both binary and multi‑class modeling. Balouchzahi et al. [9] used Twitter to collect hope‑related instances, gathering roughly 100,000 English tweets posted between January and June 2022, a period dominated by discussions of women’s reproductive rights, Black civil‑rights issues, religion, and mainstream politics. After discarding incomplete or duplicate texts, the pool dropped to just under 70,000 tweets. Filtering out messages with fewer than ten tokens reduced the set to fewer than 50,000, and the subsequent removal of retweets left about 23,000 unique, original tweets. These were shuffled to avoid topical clustering, and a random sample of 10,000 tweets was selected for manual annotation. A small number were eliminated during quality control, so the final annotated subset is slightly smaller than the initial sample, but still provides a substantial and balanced foundation for modelling hope phenomena in English social‑media discourse. The further details on annotation and pre-processing are provided in the dataset paper [9].
3.2 Spanish
For Spanish, we rely on the corpus introduced by Sidorov et al. [40] for both binary and multi‑class hope detection. The authors translated the PolyHope English trigger words into Spanish (e.g., hope → esperanza, wish → ojalá, believe → creo, dream → sueño) and asked native speakers to check that the list covered all common variants. Using these cues they first collected 82,725 keyword‑matching tweets, then added the 50,000 most recent Spanish tweets available in March 2022, yielding 132,725 raw items. After standard preprocessing—de‑identifying user mentions and URLs, removing duplicates and retweets, and discarding messages with fewer than five tokens or only emojis—the dataset was reduced to roughly 33,300 unique, content‑rich tweets. After pre-processing, only 35,000 tweets were left, and the remaining were excluded during the quality control and annotation phase. The further details on annotation and pre-processing are provided in the dataset paper [40].
3.3 Sarcasm for English and Spanish
We expanded the English part of the dataset using two benchmark datasets in sarcasm detection iSarcasmEval [41] and Sarcasm Corpus V2 [42]. The dataset is aligned with our task as they are also collected using Twitter (Now X). We combined these datasets and used our Hope classifier (the best transformer model described in [9] fine-tuned on the PolyHope dataset), filtered all instances of sarcasm that can be detected as Hope in any category. Similarly, we developed a deep learning classifier (the best transformer model described in [9] fine-tuned on the Sarcasm dataset) and filtered all instances of potential sarcasm in other classes of Hope, i.e., unrealistic hope.
The Spanish instances of Sarcasm were synthetically generated using GPT-4o, by translating the English instances and keeping the sarcastic tone in Spanish. The new labels were then verified using two expert annotators. In the binary setting we merged every tweet annotated as “sarcasm” with the broader “not‑hope” category, producing a clean dichotomy between hopeful and non‑hopeful discourse. This decision reflects the pragmatic observation that sarcastic posts rarely convey genuine hope; rather, they typically critique or ridicule the prospect of a positive outcome. The grouping of sarcasm with no hope, therefore, sharpens the semantic boundary the classifier must learn, reduces class imbalance, and minimizes the risk that ironic language will be misinterpreted as an expression of hope.
For the multi‑class task, however, we preserved sarcasm as its own label alongside hope, not‑hope, and the other original classes provided by PolyHope. Retaining sarcasm allows the model to capture the distinctive lexical and pragmatic cues of ironic speech, which differ markedly from both sincerely hopeful and straightforwardly pessimistic language. Keeping it separate also facilitates finer‑grained downstream analyses—researchers can quantify how sarcastic framing interacts with sensitive topics such as reproductive rights or racial justice—while still leveraging the same annotated tweet set described in [9] and [40]. The examples of the data, in English and Spanish, can be seen in Table 1 and 2 respectively, while Table 3 shows the statistics of the dataset.
Table 1: Sample English texts from new proposed version of PolyHope V2 with Sarcasm.
PolyHope
Text Label
#USER# You just can’t get out of your own way, can you? Not Hope
Congratulations #USER# bro for completing, my allah god bless u ahead
a good and bright future inshallah… #URL#
Generalized Hope
#USER# damn man i’m here just tryna be a nice friend and dis is how u treat me
i hope u choke with cold spaghetti
Unrealistic Hope
It was really hard exam but fortunately I studied a little last night and
I expect to pass the test
Realistic Hope
Oh, I’m sure this meeting will solve all our problems!
Sarcasm
Table 2: Sample Spanish tweets from PolyHope V2.
PolyHope-Spanish
Text Label
#USER# Que gente de m…sino saben no hablen.
Tengo mel… haya dado la teta a la melliza equivocada, entr el sueño y cans Not Hope
#USER# Uy! Ojalá. Pero tiene que ser más como el
Fight for NY porque el Icon era asquerosamente horrible. Generalized Hope
#USER# Aquí en la Mora y Piedra Azul seguimos esperando
la segunda fase de alumbrado público de sus calles #USER# #USER# Realistic Hope
Todo el maldito día trabajando en una presentación.
Ansío el día en que deje de ser una corporate slave. Unrealistic Hope
Si el sonido de la lluvia afuera no fuera ya suficientemente
malo… sería el fondo perfecto para mí mientras intento
estudiar. ¡Maravilloso! Sarcasm
Table 3: Statistics of the PolyHope V2 dataset.
Task Labels English Spanish
Hope 4,434 9,654
Binary Not Hope 5,081 10,788
Not Hope 4,081 9,788
Sarcasm 1,259 1259
Generalized Hope 2,335 5,007
Realistic Hope 982 2024
Multiclass Unrealistic Hope 858 2364
Total 9,515 20,442

### v2

@@ Abstract
Your abstract.
@@ 1 Introduction
Advances in Natural Language Processing (NLP) for emotion recognition have been significant, yet identifying the covert and complex emotion of hope remains a major challenge. Recent improvements in NLP have enhanced applications in sentiment analysis, mental health assessments, social media monitoring, and educational platforms [poria2016deeper, cambria2013new, bustos2021emotion, gomez2023emotion, mishra2025unveiling]. Despite recent progress, a persistent challenge in emotion recognition lies in identifying subtle and complex emotions, particularly hope, which is often overlooked in standard emotion taxonomies [mohammad2021sentiment]. However, these strides have not fully addressed the intricacies of detecting hope. Unlike overt emotions such as anger, fear, sadness or joy, hope is an inherently ambiguous emotion that exhibits differently across various contexts, and carries multiple emotional dimensions [buechel2022emobank, chatterjee2019semeval]. Therefore, traditional NLP techniques struggle to capture subtle emotions and highlight the need for more refined, context-aware approaches to accurately detect nuanced emotional expressions, such as hope [webber2020proceedings].
@@
Accurate detection of hope speech unlocks substantial benefits across real-world NLP applications, such as enhancing student engagement and success [xie2019trends], understanding patients’ mental states [resnik2015beyond], and mental disorder understanding (NPMDU). For example, within online education, instructors can leverage these systems as personalized support tools to accurately identify expressions of hope to better understand students’ motivations, interests, and emotional barriers—ultimately enabling targeted interventions to [xie2019trends]. Similarly, mental health professionals can leverage the natural progression of hope speech to better understand patients’ mental states, therapeutic outcomes, and concealed psychological factors over time. Research emphasizes that accurately interpreting psychological states requires identifying positive emotional expressions, particularly hope and optimism [resnik2015beyond]. As a result, NLP systems significantly benefit from incorporating hope detection technologies, since integrating positive emotional indicators alongside negative sentiments enables more comprehensive, nuanced, and effective analytic models.
@@
Traditional NLP techniques struggle to detect hope due to its complex, dual nature as defined by Lazarus [lazarus1991emotion]. According to Lazarus, hope consists of two essential components: (i) a specific desire for positive outcomes, and (ii) an overall optimistic view of future events. The inherent contradictions within hope result in complicated textual interpretations, causing individual textual markers to be variably perceived as genuine goals, imaginary objectives, or sardonic remarks. Consequently, expressions intended to reflect hopeful thinking may inadvertently suggest agitation, depending on their textual context. For example, the sentence “I hope things finally change around here” reveals optimism but also hints at frustration, depending on how it is used in the text. Most emotion recognition systems trained on simplified affective categories (e.g., Ekman’s six basic emotions) tend to fail when processing hope, as a generic positive sentiment or as non-committal discourse, failing to grasp its subtleties [ekman1992argument, buechel2022emobank]. Thus, emotion recognition systems trained on basic emotion sets fail to correctly identify hope and often misclassify it either as a simple positive emotion or as a neutral, non-committal state.
@@
Additionally, recognizing hope is further complicated by sarcasm, as sarcastic expressions frequently employ positive language to communicate negative sentiments. The mismatch between literal meanings and intended emotions makes it challenging for machine learning and deep learning classifiers to distinguish genuinely hopeful statements from their sarcastic counterparts. For instance, sarcasm occurs when someone says, “Great, I really hope this mess gets worse,” even though the words themselves convey hope, but the semantic context reveals sarcastic negativity [riloff2013sarcasm].
@@
The semantic incongruity between lexical content and emotional meaning makes emotion detection more challenging. The combination of figurative language and sarcasm complicates consistent emotion classification for computers since they primarily work with superficial word analysis instead of contextual interpretation [poria2016deeper]. Accurate context-aware emotion detection primarily depends on extensive, diverse datasets and sophisticated models [cambria2017sentiment].
However, there is a notable lack of annotated datasets explicitly related to hope, and existing datasets often either entirely omit hope or classify it merely as a subset of general positive emotions that cause incorrect predictions. Therefore, more comprehensive annotated datasets are necessary that explicitly capture hope through genuinely positive sentiments, unfounded expectations, and sarcastic expressions.
@@
In this study, we bridge the gap between emotion recognition and pragmatic language understanding and present a bilingual annotated hope speech dataset designed for hope detection in both English and Spanish language. We explicitly annotate sarcasm within hope expressions and create rich annotation guidelines for different hope categories. We design four new categories (i) generalized optimism, (ii) realistic expectations, (iii) unrealistic or wishful thinking, and (iv) sarcastic hope. To ensure accuracy and consistency in annotations, we rigorously adhered to detailed taxonomical guidelines and employed inter-annotator agreement protocols. We develop context-aware, sarcasm-resilient emotion classification models as a baseline to detect and evaluate the performance of the hope detection system in English and Spanish languages. This dataset can be used for advancing cross-linguistic emotion recognition research.
@@ 2 Literature Review
Natural Language Processing (NLP) has advanced beyond basic sentiment classification (positive, negative, or neutral) toward identifying more nuanced emotional states, including the subtle and underexplored emotion of hope [cambria2017sentiment, mohammad2021sentiment]. While emotions such as joy, fear, and regret have received substantial attention, hope remains comparatively underresearched, despite its significant role in mental health support [resnik2015beyond], education [xie2019trends], online communication, and social programs [ullah2024hope]. Hope consists of two distinct components that contribute to its complexity: (i) a general optimism inherent in individuals, and (ii) a specific expectation for positive outcomes [lazarus1991emotion]. Given its unique semantic, emotional, and contextual characteristics, recent research in hope speech detection has increasingly treated hope as an independent emotional category rather than merely grouping it under general positive emotions.
@@ 2.1 Methodologies for Hope Speech Detection
Different methodologies have been proposed for detecting hope speech and addressing challenges related to context, semantics, and multilingual content. Various studies have explored techniques to identify hope expressions within social media and textual documents [strapparava2008learning, palakodety2020hope]. For instance, Balouchzahi et al. [balouchzahi2023polyhope] introduced “PolyHope,” an NLP system that operates at semantic and contextual levels to detect nuanced expressions of hope. Similarly, another study [sidorov2023regret] leveraged transformer-based models to differentiate between expressions of hope and regret, and highlighted complexities in jointly modeling these emotionally similar pairs. Despite these advancements, current methodologies lack in distinguishing hope expressions from sarcastic or superficially hopeful messages, particularly on social media platforms (e.g., X (previously Twitter), and Reddit). Deep learning models provide poor results because they often misinterpret pragmatic signals and contextual discourse markers [taboada2016sentiment]. These limitations underline the need to incorporate pragmatic and contextual awareness to improve model accuracy. Moreover, addressing these contextual challenges becomes even more critical when extending hope detection methodologies across multilingual and cross-cultural datasets.
@@
A growing body of work has explored multilingual and cross-lingual hope speech detection, recognizing its importance in linguistically diverse societies. Chakravarthi et al. [chakravarthi2020hopeedi, chakravarthi2022overview] introduced multilingual approaches to detect hope speech in English, Tamil, Malayalam, and Kannada. The authors explicitly emphasized equality and diversity and inclusion (EDI) to detect hope speech. Nevertheless, their models did not sufficiently account for distortions caused by sarcasm or unrealistic expectations embedded within hopeful expressions. Similarly, researchers such as Arunima et al. [arunima2021ssn_dibertsity] have contributed to shared tasks like Hope Speech Detection for Equality, Diversity, and Inclusion (HopeEDI), which expanded datasets and evaluation benchmarks in multiple Indian languages. Additional efforts have targeted Urdu datasets, with studies exploring the transferability of hope speech models across languages and the need for language-specific semantic resources [balouchzahi2025urduhope, ponnusamy2024vel]. Chakravarti [chakravarthi2022multilingual] addressed hope speech detection in Tamil, English, and Malayalam, highlighting classification challenges due to code-mixed data. Moreover, Malik et al. [malik2023multilingual] explored a joint multilingual and translation-based approach for hope speech detection in English and Russian using a fine-tuned Russian-RoBERTa model, achieving 94% accuracy and 80.24However, most of these models do not explicitly address the confounding effects of sarcasm, exaggeration, or unrealistic optimism expressed through hopeful language—factors that often distort detection accuracy in real-world applications. Moreover, research on low-resource and code-mixed language scenarios remains limited, although recent work has begun to tackle these challenges through cross-lingual transfer learning and zero-shot classification approaches [thangaraj2024cross].
@@ 2.2 Existing Datasets for Hope Speech Detection
Several datasets have been introduced to facilitate advancements in hope speech detection in various languages, such as Spanish [garcia2023hope, garcia2024overview], Bengali [nath2023bonghope]. For example, the “HopeEDI” dataset [chakravarthi2020hopeedi] introduced a multilingual corpus specifically targeting hopeful expressions related to equality, diversity, and inclusion (EDI). Additionally, the “IberLEF” Hope datasets [garcia2023hope, garcia2024overview] extended hope speech detection into Spanish, and emphasized annotations around social causes and community identities. Another dataset, “BongHope” [nath2023bonghope], presented a Bengali corpus for hope speech detection. Furthermore, the HOPE shared tasks at IberLEF provided comprehensive benchmarks and detailed insights into effective methodologies for annotating hope datasets [garcia2024overview, jimenez2023overview]. In parallel, the HASOC-Hope subtrack, part of the FIRE shared tasks, introduced additional multilingual datasets covering Hindi, Malayalam, and Tamil, further enriching the research landscape in this area [mandl2020overview]). Finally, the LREC Hope Speech dataset offers a multilingual resource tailored for cross-lingual and zero-shot hope speech detection, with a special focus on low-resource language scenarios [pelicon2021zero].
@@ 2.3 Sarcasm in Hope Speech: A Complex Challenge
The detection of hope speech is highly impeded by the presence of sarcasm. Positive sentimental language through rhetorical expression functions as sarcasm to convey negative meanings while possessing a humorous or mocking quality. For example, “This one needs to go well as it happened previously.” This hopeful-sounding statement contains semantic skepticism through ironic speech. The model detection of hopeful speech can fail due to sarcastic statements that resemble hopeful statements. The improper evaluation of sentiment features occurs when basic analysis methods, such as positive polarity measurements, operate independently from detailed contextual analysis.
@@
The task of sarcasm detection is challenging because it requires contextual information, tonal cues, and background understanding [cambria2017sentiment]. Research on sarcasm detection within NLP has approached this problem in four ways: (i) Contextual embeddings (e.g., BERT, RoBERTa), (ii) Multitask learning frameworks (to jointly learn emotion and sarcasm), (iii) Contrastive learning approaches, and (iv) Discourse-aware models that incorporate surrounding sentences or conversational history. Yet sarcasm-aware hope detection remains an unexplored intersection, and no existing hope detection dataset provides labeled sarcastic examples, as emphasized by Balouchzahi et al. [balouchzahi2023polyhope] in their analysis of NPMDU-related textual data. To address this challenge, we annotate hope expressions along with sarcasm indicators and provide a richer, more realistic representation of hope in social discourse. Such annotation is critical for building models that can distinguish between sincere emotional optimism and sarcastic cynicism, as such distinction is particularly important in domains like education, counseling, and mental health.